A person is pushing a car down a horizontal road. She exerts a force of 4000 N at an angle of 10°. If the total mass of the car is 1000 kg and the µ between the car and the road is -0.35, how much work is done if the car travels 100 meters?


1
Expert's answer
2021-05-17T16:25:06-0400
"W=W_{push}-W_{fr},""W=F_{push}dcos\theta-\mu mgd,""W=4000\ N\cdot100\ m\cdot cos10^{\circ}-0.35\cdot1000\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot100\ m,""W=50923\ J."